OpenStudy (gabylovesyou):

Choose the equation of the line passing through the point (-2, 6) and parallel to y = 1/2x - 8.

OpenStudy (gabylovesyou):

y = 1/2x + 7 y = 1/2x + 5 y = 2x + 7 y = 2x + 5

OpenStudy (ghazi):

y=1/2x+7

OpenStudy (gabylovesyou):

Thank you!! :)

OpenStudy (ghazi):

:)